Solve the equation.
logbase6(x − 5) + logbase6(x) = 2
I know you can bring them together to get logbase6(x(x-5))=2
but then I'm stuck..

Answer by stanbon(75887)

Solve the equation.
logbase6(x − 5) + logbase6(x) = 2
---------
log6[x(x-5)) = 2
--------
log = EXPONENT = 2
base = 6
result = x^2-5x
------
x^2-5x = 6^2
-----
x^2-5x-36 = 0
------
(x-9)(x+4) = 0
Positive solution::
x = 9
---------------
Cheers,
Stan H.
------------------